# Corresponding Author

### v3

@@ Abstract
Generative Models, Image Synthesis, Diffusion Models, Personalization
@@ 1 Introduction
Recently, generative models have shown remarkable progress in the field of natural language processing and computer vision. Notable examples, such as ChatGPT wu2023brief and text-to-image diffusion models croitoru2023diffusion , have showcased impressive capabilities in content creation. However, these advanced models often struggle to fulfill specific requirements, such as answering domain-specific queries or accurately depicting user’s portraits. This limitation highlights the importance of Personalized Content Synthesis (PCS), which enables users to customize models for their unique tasks and requirements. As a critical area of research, PCS is increasingly recognized as essential for achieving Artificial General Intelligence (AGI), evidenced by the growing number of companies releasing products that support personalized content creation, like utilizing Reinforcement Learning to fine-tune language models uc2023survey . In this paper, we focus on PCS within the context of diffusion models in computer vision. The objective of PCS is to learn the subject of interest (SoI) from a small set of user-uploaded samples and generate images that align with user-defined contexts. For instance, as illustrated in Fig. 1, PCS can customize a unique cat into various scenarios, such as “wearing pink sunglasses” and “in the snow”.
@@
Figure 1: Given a few reference images of a subject (e.g., a cat 2ruiz2023dreambooth or face 54xiao2023fastcomposer ), PCS aims to generate new renditions of the subject that align with user-defined textual prompts. The task requires preserving the subject’s identity (fidelity) while adapting to diverse contexts. The examples are generated by this survey using DreamBooth 2ruiz2023dreambooth and InstantID 37wang2024instantid .
@@
The emergence of diffusion models has significantly facilitated text-guided content generation, leading to a rapid expansion of PCS. As depicted in Fig. 2, the number of research papers on PCS has surged over time. Following the release of key innovations such as DreamBooth 2ruiz2023dreambooth and Textual Inversion 1gal2022image in August 2022, over 150 methods have been proposed in a remarkably short period. These methods can be classified using several criteria. First, in terms of training strategy, we differentiate between test-time fine-tuning (TTF) and pre-trained adaptation (PTA) approaches. TTF methods fine-tune generative models for each personalization request during inference phase, while PTA methods strive to pre-train a unified model capable of generating a wide range of subjects of interest (SoI). Second, we categorize the techniques employed in PCS into four primary areas: attention-based operations, mask-guided generation, data augmentation, and regularization. As shown in Fig. 2, we summarize various influential PCS methods based on these criteria. Third, the scope of personalization has expanded considerably. The SoI now includes not only general objects but also extends to human faces, artistic styles, actions, and other intricate semantic elements. This growing interest encompasses the creation of complex compositions and the combination of several conditions. Additionally, the research landscape has broadened from static images to other modalities such as video and 3D representations.
@@
Figure 3: The trade-off between text alignment and visual fidelity in personalized image synthesis, illustrated through DreamBooth-generated 2ruiz2023dreambooth examples of a customized cat wearing sunglasses. Overfitting occurs when the model focuses solely on reconstructing the cat, disregarding the sunglasses context. Underfitting, on the other hand, reflects the model’s attempt to satisfy the text prompt but fails to accurately represent the personalized cat. Collapse signifies a failure to meet both criteria.
This survey aims to provide a comprehensive overview of the existing methods, frameworks, and challenges of PCS to help readers understand current trends in this evolving landscape and promote further improvements in this area. The structure of this survey is organized into several key sections: In Sec. 2, We begin with a brief introduction to diffusion models for better understanding of PCS. In Sec. 3, we introduce two primary frameworks for PCS: test-time fine-tuning and pre-trained adaptation. In Sec. 4, we categorize common techniques into four main parts, attention-based methods, mask-based generation, data augmentation, and regularization. In Sec. 5, we summarize various PCS methods based on different personalization tasks, including personalization on object, face, style, and etc. In Sec. 6, we demonstrate PCS’s growing impact in video, 3D, and other areas beyond traditional image generation tasks. In Sec. 7, we review the current evaluation dataset and metrics and construct a benchmark on the existing methods for further research. In Sec. 8, we discuss current challenges faced in PCS and propose potential future directions to overcome these problems.
@@ 2 Fundamentals
Modern diffusion modeling has evolved into a sophisticated framework that unifies discrete and continuous generative paradigms through stochastic differential equations (SDEs) song2020score and ordinary differential equations (ODEs) song2020score ; karras2022elucidating . This section introduces the core mathematical foundations, emphasizing innovations in the diffusion process and conditional mechanisms critical for personalized generation.
@@ 2.1 Denoising Diffusion Probabilistic Models (DDPMs)
The architecture of diffusion models comprises two complementary processes: a forward diffusion that systematically perturbs data distributions, and a learned reverse process that reconstructs signals through iterative refinement. Originating from discrete-time Markov chains in Denoising Diffusion Probabilistic Model (DDPM) ho2020denoising , the forward process applies Gaussian noise corruption over $T$ steps:
@@
$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),\beta_{t}I\right)$ (2)
@@ 2.2 Stochastic Differential Equations (SDEs)
To overcome these limitations, the framework was generalized through continuous-time SDE song2020score , which provide a unified perspective for modeling diffusion dynamics:
@@
where $\mathbf{f}(\mathbf{x},t)$ encodes deterministic drift components, $g(t)$ modulates stochastic diffusion via Wiener process $\mathbf{w}$. This continuous perspective subsumes DDPMs as special cases while enabling adaptive noise scheduling strategies such as variance-preserving (VP) and variance-exploding (VE) schedules through strategic choices of $\mathbf{f}$ and $g$ karras2022elucidating .
For the reverse process, a crucial result anderson1982reverse shows that the reverse diffusion process also follows an SDE:
@@ 2.3 Ordinary Differential Equations (ODEs)
These limitations motivate the derivation of deterministic sampling trajectories through probability flow ordinary differential equations (ODEs) song2020score ; karras2022elucidating , obtained by eliminating the stochastic term from the SDE formulation:
@@
The ODE’s deterministic nature arises from two synergistic components: the original drift term $\mathbf{f}(\mathbf{x},t)$ and a correction term involving the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. This score-guided adjustment preserves the marginal data distribution $p_{t}(\mathbf{x})$ while eliminating pathwise stochasticity. Crucially, the resulting trajectories exhibit intrinsic geometric regularity, allowing the complex denoising process to operate in a simplified mathematical space. This enabling 5-10x faster sampling than SDEs through adaptive ODE solvers like DPM-Solver lu2022dpm and DPM-Solver++ lu2022dpmplus . Recent work chen2024trajectory further identifies strong geometric regularity in these ODE-based sampling trajectories, demonstrating that they inherently follow a linear-nonlinear-linear structure regardless of generated content. This trajectory regularity enables dynamic programming-based time scheduling optimization with negligible computational overhead.
@@ 2.4 Conditional Generation Mechanisms
where the conditioning signal $c$ can be integrated via multiple synergistic mechanisms, such as cross-modal attention rombach2022high and spatial modulation zhang2023adding .
This conditional paradigm directly enables the applications of PCS. Leading text-to-image systems, such as Stable Diffusion (SD) rombach2022high and DALLE ramesh2022hierarchical , are widely adopted to empower users in controlling customized content through text instructions.
@@ 3.1 Test-time Fine-tuning (TTF) Framework
TTF methods represent the foundational approach for personalized synthesis, where models adapt to new subjects through instance-specific optimization during inference. As illustrated in Fig. 4, this framework operates through two core principles: 1) Test-time adaptation fine-tunes model parameters to learn the key visual element of the SoI. 2) Semantic-aware modifier system represents the SoI at the token level to bridge the gap between visual adaptation and textual control. In the following Sec. 3.1.1 and Sec. 3.1.2, we will detail these principles respectively.
@@ 3.1.1 Test-time Adaption
where $c$ represents the condition signal, typically the reference image caption. Compared to the large-scale pre-training described in Eq. 7, key differences lie in the training data and learnable parameters. The training samples are typically restricted to the SoI references, sometimes supplemented with a regularization dataset to mitigate overfitting 2ruiz2023dreambooth . For the selection of learnable parameters $\mathbf{\theta^{\prime}}$, the commonly adopted options include token embeddings 1gal2022image ; 23voynov2023p+ , the entire diffusion model 2ruiz2023dreambooth ; 151zhang2024generative , specific subsets of parameters 6kumari2023multi ; 7tewel2023key ; 8han2023svdiff , or introducing new parameters such as adapters 72xiang2023closer ; 5sohn2024styledrop , and LoRA 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg , which will be discussed in Sec. 3.1.3.
Test-time generation. The test-time generation is proceeded once the model has been fine-tuned with the optimized parameters $\mathbf{\theta^{\prime}}$. By composing novel input prompts that incorporate the SoI’s unique identifier (introduced in Sec. 3.1.2), the adapted model can synthesize diverse images while preserving the subject’s distinctive characteristics. This approach enables flexible control over the generated content through natural language descriptions.
@@ 3.1.2 Unique Modifier
The unique modifier is a critical component in PCS, enabling text-based representation of the SoI for flexible prompt composition. As illustrated in Fig. 4, this modifier serves as a textual description for the SoI, allowing combination with other descriptions (e.g., “V* on the beach”) during inference. Normally, the construction of the unique modifier can be divided into three categories:
Learnable token embedding 1gal2022image ; 23voynov2023p+ ; 61alaluf2023neural ; 30zhang2023prospect ; 48wang2023hifi ; 66pang2023cross . This method adds a new token and its corresponding embedding vector to the word dictionary of the tokenizer. An intuitive example is that this method creates a new word that does not exist in the dictionary. This word acts as the modifier, with adjustable weights during fine-tuning, while the embeddings of other tokens in the pre-defined dictionary will not be influenced. This approach requires only a few kilobytes of additional parameters and maintains the base model’s capabilities for non-customized generation.
Plain text 2ruiz2023dreambooth ; 6kumari2023multi ; 7tewel2023key ; 5sohn2024styledrop ; 151zhang2024generative . This approach utilizes an explicit text description to represent the SoI. For example, words such as cat could directly represent the user’s cat in the references. This setting typically requires fine-tuning the parameters of the diffusion model components such as UNet ronneberger2015unet or Transformer blocks vaswani2017attention . The plain text offers user-friendly prompt construction and injects subject prior information to ease tuning difficulty. However, this technique may over-specialize common terms, limiting their broader applicability as the model learns to associate general words with specific SoI characteristics.
Rare token 2ruiz2023dreambooth . This technique employs infrequently used tokens to minimize their impact on commonly used vocabulary. Similar to the plain text approach, the embeddings of rare tokens remain unchanged during fine-tuning. However, these rare tokens fail to provide useful subject prior information and still exhibit weak interference with unrelated vocabulary, potentially leading to ambiguity between the original meaning and the intended SoI reference.
@@ 3.1.3 Training Parameter Selection
The selection of trainable parameters represents a critical design consideration in PCS, directly impacting many key performance metrics: subject fidelity, training efficiency, and model storage requirements. Current tunable parameters can be broadly categorized into token embedding, existing model components, parameter-efficient techniques, and hybrid approaches.
Token embedding. As introduced in Sec. 3.1.2, token embedding optimization 1gal2022image introduces learnable tokens as the unique modifier to represent SoI through noise-to-image reconstruction. While achieving remarkable parameter efficiency, this approach faces challenges in detail preservation and prolonged training time due to the inherent compression of complex features into low-dimensional embeddings. Subsequent works 23voynov2023p+ ; 61alaluf2023neural ; 30zhang2023prospect ; 48wang2023hifi ; 66pang2023cross aim to address these limitations through different strategies, which are summarized in Sec. 5.1.
Existing model parameters. This paradigm directly optimizes pre-trained components, such as the text encoder, U-Net blocks, and transformer layers 2ruiz2023dreambooth ; 151zhang2024generative ; 6kumari2023multi ; 7tewel2023key ; 8han2023svdiff . Benefitting from the advanced representation capacity of these modules, the fine-tuning phase can achieve faster convergence (5-10 minutes) and superior visual fidelity compared to token-only methods, though at the cost of significant storage overhead. In addition, these modules inherently support attention mechanisms to facilitate feature enhancement operations.
Parameter-efficient extensions. Recent advanced methods have introduced parameter-efficient techniques into PCS, such as LoRA 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg ; 51lu2024object ; 95achlioptas2023stellar ; 76yeh2023navigating and adapter modules 72xiang2023closer ; 5sohn2024styledrop , which inject small, trainable components into the base model. These methods achieve comparable performance to full parameter fine-tuning while dramatically reducing storage requirements.
Combined strategy. Since the aforementioned strategies are not conflicting, some methods allocate different learning rates and training phases to each component type to achieve optimal balance between fidelity and efficiency. For instance, the fine-tuned token embedding can be regarded as an effective initialization for the subsequent model weight fine-tuning 90hyung2023magicapture . Also, these two parts can be simultaneously optimized with different learning rates 74avrahami2023break ; 86hua2023dreamtuner .

### v4

@@ Abstract
Generative Models, Image Synthesis, Diffusion Models, Personalized Content Synthesis, Subject Customization
00footnotetext: Accepted by Machine Intelligence Research, DOI: 10.1007/s11633-025-1563-3
@@ 1 Introduction
Recently, generative models have shown remarkable progress in the field of natural language processing and computer vision. Notable examples, such as ChatGPT [1] and text-to-image diffusion models [2], have showcased impressive capabilities in content creation. However, these advanced models often struggle to fulfill specific requirements, such as answering domain-specific queries or accurately depicting user’s portraits. This limitation highlights the importance of Personalized Content Synthesis (PCS), which enables users to customize models for their unique tasks and requirements. As a critical area of research, PCS is increasingly recognized as essential for achieving Artificial General Intelligence (AGI), evidenced by the growing number of companies releasing products that support personalized content creation, like utilizing Reinforcement Learning to fine-tune language models [3]. In this paper, we focus on PCS within the context of diffusion models in computer vision. The objective of PCS is to learn the subject of interest (SoI) from a small set of user-uploaded samples and generate images that align with user-defined contexts. For instance, as illustrated in Fig. 1, PCS can customize a unique cat into various scenarios, such as “wearing pink sunglasses” and “in the snow”.
@@
Figure 1: Given a few reference images of a subject (e.g., a cat [4] or face [5]), PCS aims to generate new renditions of the subject that align with user-defined textual prompts. The task requires preserving the subject’s identity while adapting to diverse contexts. The examples are generated by this survey using DreamBooth [4] and InstantID [6].
@@
The emergence of diffusion models has significantly facilitated text-guided content generation, leading to a rapid expansion of PCS. As depicted in Fig. 2, the number of research papers on PCS has surged over time. Following the release of key innovations such as DreamBooth [4] and Textual Inversion [7] in August 2022, over 150 methods have been proposed in a remarkably short period. These methods can be classified using several criteria. First, in terms of training strategy, we differentiate between test-time fine-tuning (TTF) and pre-trained adaptation (PTA) approaches. TTF methods fine-tune generative models for each personalization request during inference phase, while PTA methods strive to pre-train a unified model capable of generating a wide range of subjects of interest (SoI). Second, we categorize the techniques employed in PCS into four primary areas: attention-based operations, mask-guided generation, data augmentation, and regularization. As shown in Fig. 2, we summarize various influential PCS methods based on these criteria. Third, the scope of personalization has expanded considerably. The SoI now includes not only general objects but also extends to human faces, artistic styles, actions, and other intricate semantic elements. This growing interest encompasses the creation of complex compositions and the combination of several conditions. Additionally, the research landscape has broadened from static images to other modalities such as video and 3D representations.
@@
Figure 3: The trade-off between text alignment and visual fidelity in personalized image synthesis, illustrated through DreamBooth-generated [4] examples of a customized cat wearing sunglasses. Overfitting occurs when the model focuses solely on reconstructing the cat, disregarding the sunglasses context. Underfitting, on the other hand, reflects the model’s attempt to satisfy the text prompt but fails to accurately represent the personalized cat. Collapse signifies a failure to meet both criteria.
@@
This survey aims to provide a comprehensive overview of the existing methods, frameworks, and challenges of PCS to help readers understand current trends in this evolving landscape and promote further improvements in this area. The structure of this survey is organized into several key sections: In Section 2, We begin with a brief introduction to diffusion models for better understanding of PCS. In Section 3, we introduce two primary frameworks for PCS: test-time fine-tuning and pre-trained adaptation. In Section 4, we categorize common techniques into four main parts, attention-based methods, mask-based generation, data augmentation, and regularization. In Section 5, we summarize various PCS methods based on different image personalization tasks, including personalization on object, face, style, and etc. In Section 6, we demonstrate PCS’s growing impact in video, 3D, and other areas beyond traditional image generation tasks. In Section 7, we review the current evaluation dataset and metrics, and introduce a benchmark on the existing methods to promote further research. In Section 8, we discuss current challenges faced in PCS and propose potential future directions in this area. Additionally, we present a comprehensive summary of all relerated papers in Table 2, Table 3, and Table 4.
@@ 2 Fundamentals
Modern diffusion modeling has evolved into a sophisticated framework that unifies discrete and continuous generative paradigms through stochastic differential equations (SDEs) [8], ordinary differential equations (ODEs) [8, 9]. This section introduces the core mathematical foundations, emphasizing innovations in the diffusion process and conditional mechanisms. By covering these theoretical developments, this section lays the groundwork for the personalization tasks explored in subsequent sections.
@@ 2.1 Denoising Diffusion Probabilistic Models (DDPMs)
The architecture of diffusion models comprises two complementary processes: a forward diffusion that systematically perturbs data distributions, and a learned reverse process that reconstructs signals through iterative refinement. Originating from discrete-time Markov chains in Denoising Diffusion Probabilistic Model (DDPM) [10], the forward process applies Gaussian noise corruption over $T$ steps:
@@
$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right)=\\ \mathcal{N}\left(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),\beta_{t}I\right)$ (2)
@@
Once training is finished, the model can apply the learned reverse diffusion transitions to any arbitrary noise input, progressively denoising it into a coherent data sample. This capability allows DDPMs to function as general-purpose generative engines, producing high-quality outputs from random noise vectors.
@@ 2.2 Stochastic Differential Equations (SDEs)
To overcome these limitations, the framework was generalized through continuous-time SDE [8], which provide a unified perspective for modeling diffusion dynamics:
@@
where $\mathbf{f}(\mathbf{x},t)$ encodes deterministic drift components, $g(t)$ modulates stochastic diffusion via Wiener process $\mathbf{w}$. This continuous perspective subsumes DDPMs as special cases while enabling adaptive noise scheduling strategies such as variance-preserving (VP) and variance-exploding (VE) schedules through strategic choices of $\mathbf{f}$ and $g$ [9].
For the reverse process, a crucial result [11] shows that the reverse diffusion process also follows an SDE:
@@ 2.3 Ordinary Differential Equations (ODEs)
These limitations motivate the derivation of deterministic sampling trajectories through probability flow ordinary differential equations (ODEs) [8, 9], obtained by eliminating the stochastic term from the SDE formulation:
@@
The ODE’s deterministic nature arises from two synergistic components: the original drift term $\mathbf{f}(\mathbf{x},t)$ and a correction term involving the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. This score-guided adjustment preserves the marginal data distribution $p_{t}(\mathbf{x})$ while eliminating pathwise stochasticity. Crucially, the resulting trajectories exhibit intrinsic geometric regularity, allowing the complex denoising process to operate in a simplified mathematical space. This enabling 5-10x faster sampling than SDEs through adaptive ODE solvers like DPM-Solver [12] and DPM-Solver++ [13]. Recent work [14] further identifies strong geometric regularity in these ODE-based sampling trajectories, demonstrating that they inherently follow a linear-nonlinear-linear structure regardless of generated content. This trajectory regularity enables dynamic programming-based time scheduling optimization with negligible computational overhead.
@@ 2.4 Conditional Generation Mechanisms
where the conditioning signal $c$ can be integrated via multiple synergistic mechanisms, such as cross-modal attention [15] and spatial modulation [16].
This conditional paradigm directly enables the applications of PCS. Leading text-to-image systems, such as Stable Diffusion (SD) [15] and DALLE [17], are widely adopted to empower users in controlling customized content through text instructions.
@@ 3.1 Test-time Fine-tuning (TTF) Framework
TTF methods represent the foundational approach for personalized synthesis, where models adapt to new subjects through instance-specific optimization during inference. As illustrated in Fig. 4, this framework operates through two core principles: 1) Test-time adaptation fine-tunes model parameters to learn the key visual element of the SoI. 2) Semantic-aware modifier system represents the SoI at the token level to bridge the gap between visual adaptation and textual control. In the following Section 3.1.1 and Section 3.1.2, we will detail these principles respectively.
@@ 3.1.1 Test-time Adaption
where $c$ represents the condition signal, typically the reference image caption. Compared to the large-scale pre-training described in Equation 7, key differences lie in the training data and learnable parameters. The training samples are typically restricted to the SoI references, sometimes supplemented with a regularization dataset to mitigate overfitting [4]. For the selection of learnable parameters $\mathbf{\theta^{\prime}}$, the commonly adopted options include token embeddings [7, 18], the entire diffusion model [4, 19], specific subsets of parameters [20, 21, 22], or introducing new parameters such as adapters [23, 24], and LoRA [25, 26, 27], which will be discussed in Section 3.1.3.
Test-time generation. The test-time generation is proceeded once the model has been fine-tuned with the optimized parameters $\mathbf{\theta^{\prime}}$. By composing novel input prompts that incorporate the SoI’s unique identifier (introduced in Section 3.1.2), the adapted model can synthesize diverse images while preserving the subject’s distinctive characteristics. This approach enables flexible control over the generated content through natural language descriptions.
@@ 3.1.2 Unique Modifier
A unique modifier is a textual token or short phrase that uniquely represents a SoI, enabling text-based representation of the SoI for flexible prompt composition. As illustrated in Fig. 4, this modifier serves as a textual description for the SoI, allowing combination with other descriptions (e.g., “V* on the beach”) during inference. Normally, the construction of the unique modifier can be divided into three categories:
Plain text [4, 20, 21, 24, 19]. This approach utilizes an explicit text description to represent the SoI. For example, words such as cat could directly represent the user’s cat in the references. This setting typically requires fine-tuning the parameters of the diffusion model components, such as UNet [28] or Transformer blocks [29], to enable the model to associate the SoI’s visual features with the plain‐text token. The plain text offers user-friendly prompt construction and injects subject prior information to ease tuning difficulty. However, this technique may over-specialize common terms, limiting their broader applicability as the model learns to associate general words with specific SoI characteristics.
Rare token [4]. This technique employs infrequently used tokens to minimize their impact on commonly used vocabulary. Similar to the plain text approach, the embeddings of rare tokens remain unchanged during fine-tuning. However, these rare tokens often fail to provide useful subject prior information and still exhibit weak interference with unrelated vocabulary, potentially leading to ambiguity between the original meaning and the intended SoI reference.
Learnable token embedding [7, 18, 30, 31, 32, 33]. This method adds a new token and its corresponding embedding vector to the dictionary of the tokenizer. An intuitive example is that this method creates a new word that does not exist in the dictionary. This inserted token has adjustable weights during fine-tuning, while the embeddings of other tokens in the pre-defined dictionary will remain unchanged. This approach requires only a few kilobytes of additional parameters and maintains the base model’s capabilities for non-customized generation. Similar to the rare token approach, it is not very user-friendly in practice, since users must learn and remember an unfamiliar token to reference their subject of interest.
@@ 3.1.3 Training Parameter Selection
The selection of trainable parameters represents a critical design consideration in PCS, directly impacting many key performance metrics: subject fidelity, training efficiency, and model storage requirements. Current tunable parameters can be broadly categorized into the following four types.
Token embedding. As introduced in Section 3.1.2, token embedding optimization [7] introduces learnable tokens as the unique modifier to represent SoI through noise-to-image reconstruction. While achieving remarkable parameter efficiency, this approach faces challenges in detail preservation and prolonged training time (typically larger than 20 minutes) due to the inherent compression of complex features into low-dimensional embeddings. Subsequent works [18, 30, 31, 32, 33] aim to address these limitations through different strategies, which are summarized in Section 5.1.
Existing model parameters. This paradigm directly optimizes pre-trained model components, such as the text encoder, U-Net blocks, and transformer layers [4, 19, 20, 21, 22]. Benefitting from the advanced representation capacity of these modules, the fine-tuning phase can achieve faster convergence (5-10 minutes) and superior visual fidelity compared to token-only methods, though at the cost of significant storage overhead. In addition, these modules inherently support attention mechanisms to facilitate feature enhancement operations.
Parameter-efficient extensions. Recent advanced methods have introduced parameter-efficient techniques into PCS, such as LoRA [25, 26, 27, 34, 35, 36] and adapter modules [23, 24], which inject small, trainable components into the base model. These methods achieve comparable performance to full parameter fine-tuning while dramatically reducing storage requirements.
Combined strategy. Since the aforementioned strategies are not conflicting, some methods allocate different learning rates and training phases to each component type to achieve optimal balance between fidelity and efficiency. For instance, the fine-tuned token embedding can be regarded as an effective initialization for the subsequent model weight fine-tuning [37]. Also, these two parts can be simultaneously optimized with different learning rates [38, 39].